OpenStudy (anonymous):

Find f if f"(theta)=sin (theta) + cos (theta), f(0)=3, and f'(0)=4

OpenStudy (anonymous):

\[f \prime \prime (\theta)=\sin \theta +\cos \theta\] \[f \prime (\theta) =-\cos \theta + \sin \theta +c\] \[f \prime (0) =-1+0+c=4\] \[c=5\] \[f \prime (\theta) =-\cos \theta + \sin \theta +5\] \[f (\theta) =-\cos \theta - \sin \theta +c\] \[f(0)=-1+c=3\] \[c=4\] f(θ)=−cosθ−sinθ+4